Suppose we experiment with an Atwood machine consisting of two masses m and 2m connected by an inextensible nonslip string running over a pulley of radius a and mass 4m

A)Determine the total kinetic energy of the system. (6marks)

B)Find the gravitational potential energy of the system. (4)

C)briefly discuss why no work is done by the forces of constrait (5)

D)what is the total energy of the system (2)

E)Find the acceleration dv/dt of the blocks


1
Expert's answer
2022-04-19T10:00:41-0400

"E=(2m-m)ga=mga,"

"g'=g\frac{2m-m}{4m}=0. 25g."